1. Calculate the cash dividends paid out to common stockholders based on the following information: Balance...

1.

Calculate the cash dividends paid out to common stockholders based on the following information:

Balance of retained earnings, December 31, 2016 =

411,000,000
Net Income for 2017       30,000,000
Dividends paid to preferred stockholders 1,000,000
Balance of retained earnings, December 31, 2017 375,000,000

2.XYZ Company has a cash balance of 7,000 and an accounts receivable balance of 2,000. Their accounts payable balance is 13,000 and the company incurred research and development expenses during the year of 3,000 which have already been paid. Their inventory balance is 3,500. Calculate the working capital. (Answer in whole $s with no comma or $ sign needed, e.g. xxxxx or 12500)

3.Webster Company has a beginning balance in Retained Earnings of $8,000. They earned a net income of $4,000 during the year and paid out $9,000 in dividends. What is the ending balance in Webster Companies Retained Earnings account at the end of the year? (Answer in whole $s without the $ sign or comma, i.e. 12345.)
Answer #1

1. Cash dividend paid out to common stockholders = Balance of retained earnings, December 31, 2016 + Net income for 2017 - Dividends paid to preferred stockholders - Balance of retained earnings, December 31, 2017 = 411,000,000 + 30,000,000 - 1,000,000 - 375,000,000 = 65,000,000

2. Current assets = Cash balance + Accounts receivables + Inventory = 7000 + 2000 + 3500 = 12500

Current liabilities = Accounts payables = 13000

Working capital = Current assets - Current liabilities = 12500 - 13000 = - 500

3. Ending balance in retained earnings = Beginning balance in retained earnings + Net income - Dividends = 8000 + 4000 - 9000 = 3000